F(x)= x^3-5x^2+144x-720
I got one of the zeroes (x=5) I need the other 2... they should be imaginary numbers.

Answer by stanbon(75887)

F(x)= x^3-5x^2+144x-720
I got one of the zeroes (x=5) I need the other 2... they should be imaginary numbers.
--------------------------------
Use synthetic division to find the quadratic factor:
---
5)....1....-5....144....-720
.......1.....0....144...|..0
------------------
Quotient: x^2+144
Remainder: 0
------
Solve x^2+144 = 0
(x-12i)(x+12i) = 0
x = 12i or x = -12i
==========================
Final Answer: x = 5, or x = 12i, or x = -12i
=================================================
Cheers,
Stan H.